# Observation of a non-Hermitian supersonic mode

### v1

@@ IV Acknowledgement
We thank Sarang Gopalakrishnan, Tim Hsieh, Dvira Segal, Nathan Wiebe, Roeland Wiersema, Cenke Xu, and Yijian Zou for their insightful discussions. We extend gratitude to Michael Foss-Feig and Peter Groszkowski for their invaluable assistance in accessing Quantinuum’s hardware resources. J.C. acknowledges support from the Shared Hierarchical Academic Research Computing Network (SHARCNET), Compute Canada, and the Canadian Institute for Advanced Research (CIFAR) AI chair program. Y.Z., J.C., and Y.B.K. were supported by the Natural Science and Engineering Research Council (NSERC) of Canada. Y.Z. and Y.B.K. acknowledge support from the Center for Quantum Materials at the University of Toronto. Y.Z. was further supported by a CQIQC fellowship at the University of Toronto, and in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725, and, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute.www.vectorinstitute.ai/#partners.
@@ A.1 Non-Hermitian physics: a motivating example
We argued in the introduction and Fig. 1 that a physical way to motivate non-Hermitian physics is from open quantum system dynamics. Usually, the Linbladian equation is used to describe the Markovian evolution of a system interacting with a thermal bath [151]:
@@
$=-i(H_{\rm eff}\rho-\rho H_{\rm eff}^{\dagger})+\sum_{i}\gamma_{i}L_{i}\rho L^{\dagger}_{i}$ (11)
@@
The last term in Eq. 11, $\sum_{i}\gamma_{i}L_{i}\rho L^{\dagger}_{i}$, describes quantum jump processes that can be associated with a measurable physical quantity, such as a spontaneous emission of a photon. Depending on the number of quantum jumps, the stochastic time evolution can be classified into different quantum trajectories. Now, if we postselect on the absence of quantum jumps, we end up with a Schrodinger-like time evolution with a non-Hermitian Hamiltonian:
@@ A.2 Sequential circuits
$|\Psi\rangle=\sum_{\{s_{x}\}_{x=1}^{n}}\ell^{T}A^{s_{1}}A^{s_{2}}\dots|s_{1}s_{2}\dots\rangle$ (13)
@@
Algorithm 3 Generating sequential quantum circuits from MPS
@@
A.3 Analytical proof of Thm. I.1
@@
In the second from last step, we use the fact that each entry of a Haar state vector consists of two standard normal variables and the weights on each string should follow the Porter-Thomas distribution: ${\rm PT}(p_{s})=2^{n}e^{2^{n}p_{s}}$, and a sum over $\operatorname{poly}(n)$ terms would converge to its mean value. We may also ignore the $p_{0^{n}}e^{-2n}$ in the denominator as it is exponentially small in $n$.
@@
Combining Lem. A.3 and Lem. A.6 completes the proof of Thm. I.1. Further, as we explain in Appx. C, each single-qubit dissipation channel can be implemented with a two-qubit gate and post-selection on the single ancillary qubit. It turns out that the result we have proved can be summarized in a quantum complexity language: With overwhelming probability, Prob. A.2 can be solved by a quantum circuit at merely constant depth when postselection is allowed (denoted as $\mathsf{PostQNC_{0}}$), but also with overwhelming probability, it cannot be solved by a polynomial-size quantum circuit. Namely,
@@
There exists an oracle $\mathcal{O}$ such that $\mathsf{PostQNC_{0}}^{\mathcal{O}}\not\subseteq\mathsf{BQP}^{\mathcal{O}}$; relative to the same oracle, $\mathsf{BQP}^{\mathcal{O}}\subset\mathsf{PostBQP}^{\mathcal{O}}$.
@@ B.1 Failure of energy minimization due to non-Hermicity
B.2 Testing with a dissipative XXZ model
@@
B.3 Initialization with warm start
@@ C.1 Quantinuum H1 Trapped ion computer
However, the high performance of the processor is accompanied by a relatively low clock rate. To mitigate the large sampling overhead of variational algorithms, we chose to perform classical optimization of circuit parameters and implemented only the optimized circuit in hardware. Nevertheless, this optimization loop could potentially be implemented on a quantum processor in the near future. We also computationally simulated the time-evolved state classically, although this could theoretically be achieved on a quantum computer using algorithms such as quantum imaginary time evolution [115]. The total experimental cost amounted to approximately 12,000 credits, with half allocated to dynamical simulation and the remainder to ground state preparation.
@@ C.2 Non-Hermitian fermionic quench
The transformed Hamiltonian is still a sum of local terms, i.e., $H=\sum_{j}h_{j,j+1}$, and one can perform the following Trotter decomposition suitable for this case [171]:
@@
For the total evolution time $t=2$ considered in Sec. I, we numerically found that a state infidelity of $0.0231-\mathcal{F}=$ 0.023 can be achieved at $m=3$, which matches the target infidelity considered by VQC.
@@ C.3 Qubit-efficient preparation of dissipative Ising ground state
In this section, we detail the resource used in conducting the experiment in Fig. 3.
@@
    Gate count. The ladder circuit (as shown in Fig. 2) at each physical site has depth $\tau$, and the gate count per site $\sim\tau q$. The gate count has a linear dependence on the system size, and the total gate count $\sim\tau qn$. To be precise, for us, $\tau=3,n=20$, and the total SU(4) gate count is 80.
@@
    Gate set. In our tensor-network algorithm, each tenor is assumed to be a SU(4) gate, which can be synthesized with three hardware-natural two qubit gates and other single qubit gates [173].

### v2

@@ VII Acknowledgement
We thank Sarang Gopalakrishnan, Tim Hsieh, Lin Lin, Dvira Segal, Nathan Wiebe, Roeland Wiersema, Cenke Xu, and Yijian Zou for their insightful discussions. We extend gratitude to Michael Foss-Feig and Peter Groszkowski for their invaluable assistance in accessing Quantinuum’s hardware resources and thank the reviewers for their valuable feedback. J.C. acknowledges support from the Shared Hierarchical Academic Research Computing Network (SHARCNET), Compute Canada, and the Canadian Institute for Advanced Research (CIFAR) AI chair program. Y.Z., J.C., and Y.B.K. were supported by the Natural Science and Engineering Research Council (NSERC) of Canada. Y.Z. and Y.B.K. acknowledge support from the Center for Quantum Materials at the University of Toronto. Y.Z. was further supported by a CQIQC fellowship at the University of Toronto, and in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP). This research used resources of the Oak Ridge Leadership Computing Facility, which is a DOE Office of Science User Facility supported under Contract DE-AC05-00OR22725, and, in part, by the Province of Ontario, the Government of Canada through CIFAR, and companies sponsoring the Vector Institute www.vectorinstitute.ai/#partners.
VIII Author Contributions
J.C., Y.B.K. and Y.Z. conceived the project and designed the theoretical framework. J.C. and Y.B.K. supervised the project and contributed crucial insights on interpreting the results. Y.Z. performed numerical simulations, executed the quantum circuits, and analyzed the results. Y.Z. wrote the manuscript with input from all authors.
IX Competing Interests
The authors declare no competing interests.
@@ A.1 Non-Hermitian physics: a motivating example
We argued in the main text that a physical way to motivate non-Hermitian physics is from open quantum system dynamics. Usually, the Linbladian equation is used to describe the Markovian evolution of a system interacting with a thermal bath [54]:
@@
$=-i(H_{\rm eff}\rho-\rho H_{\rm eff}^{\dagger})+\sum_{i}\gamma_{i}L_{i}\rho L^{\dagger}_{i}.$ (40)
@@
The last term in Eq. 40, $\sum_{i}\gamma_{i}L_{i}\rho L^{\dagger}_{i}$, describes quantum jump processes that can be associated with a measurable physical quantity, such as a spontaneous emission of a photon. The stochastic time evolution can be classified into different quantum trajectories depending on the number of quantum jumps. Now, if we postselect on the absence of quantum jumps, we end up with a Schrodinger-like time evolution with a non-Hermitian Hamiltonian:
@@ A.2 Review of sequential circuits
$|\Psi\rangle=\sum_{\{s_{x}\}_{x=1}^{n}}\ell^{T}A^{s_{1}}A^{s_{2}}\dots|s_{1}s_{2}\dots\rangle.$ (42)
@@
Box 5 Generating sequential quantum circuits from MPS
@@
A.3 Analytical proof of Thm. I.1 and implication in complexity theory
@@
In the second from last step, we use the fact that each entry of a Haar state vector consists of two standard normal variables and the weights on each string should follow the Porter-Thomas distribution: ${\rm PT}(p_{s})=2^{n}e^{-2^{n}p_{s}}$, and a sum over $\operatorname{poly}(n)$ terms would converge to its mean value. We may also ignore the $p_{0^{n}}e^{-2n}$ in the denominator as it is exponentially small in $n$.
@@
Combining Lem. A.3 and Lem. A.6 completes the proof of Thm. I.1. Further, as we explain in Appx. C, each single-qubit dissipation channel can be implemented with a two-qubit gate and post-selection on the single ancillary qubit. It turns out that the result we have proved can be summarized in a computational complexity language: With overwhelming probability, Prob. A.2 can be solved by a quantum circuit at merely constant depth when postselection is allowed (denoted as $\mathsf{PostQNC_{0}}$), but also with overwhelming probability, it cannot be solved by a polynomial-size quantum circuit. Namely,
@@
There exists a quantum oracle $\mathcal{O}$ (as defined in Prob. A.2) such that $\mathsf{PostQNC_{0}}^{\mathcal{O}}\not\subseteq\mathsf{BQP}^{\mathcal{O}}$; relative to the same oracle, $\mathsf{BQP}^{\mathcal{O}}\subset\mathsf{PostBQP}^{\mathcal{O}}$.
@@ B.1 Failure of energy minimization due to non-Hermicity
B.2 Testing the algorithm with a dissipative XXZ model
@@
B.3 The importance of warm start
@@ C.1 Quantinuum H1 Trapped ion computer
However, the high performance of the processor is accompanied by a relatively low clock rate. To mitigate the large sampling overhead of variational algorithms, we chose to perform classical optimization of circuit parameters and implemented only the optimized circuit in hardware. In principle, this optimization loop could potentially be implemented on a quantum processor in the near future. We also computationally simulated the time-evolved state classically, although this could be achieved on a quantum computer using algorithms such as quantum imaginary time evolution [79]. The total experimental cost amounted to approximately 12,000 credits, with half allocated to dynamical simulation and the remainder to ground state preparation.
@@ C.2 Non-Hermitian fermionic quench
The transformed Hamiltonian is still a sum of local terms, i.e., $H=\sum_{j}h_{j,j+1}$, and one can perform the following Trotter decomposition suitable for this case:
@@
For the total evolution time $t=2$ considered in the work, we numerically found that a state infidelity of $0.0231-\mathcal{F}=$ 0.023 can be achieved at $m=3$, which matches the target infidelity considered by VQC.
@@ C.3 Qubit-efficient preparation of dissipative Ising ground state
In this section, we detail the resources used in experimental preparation of the dissipative Ising ground state.
@@
    Gate count. The ladder circuit (as shown in Fig. 2) at each physical site has depth $\tau$, and the gate count per site $\sim\tau q$. The gate count has a linear dependence on the system size, and the total gate count $\sim\tau qn$. To be precise, for us, $q=2,\tau=2$, and $n=20$, thus the total SU(4) gate count is 80.
@@
    Gate set. In our tensor-network algorithm, each tenor is assumed to be an SU(4) gate, which can be synthesized with three hardware-natural two-qubit gates and other single-qubit gates [81].